OpenStudy (anonymous):

The equation of a line with slope and including point (-3, 5) is y = - 1/4x -4 true or false?

OpenStudy (anonymous):

we are missing part of the question, what comes after the word "slope" and before "and"

OpenStudy (anonymous):

m = - 1/4

OpenStudy (anonymous):

now plug in x = -3 and y = 5 into y = - 1/4x -4 5 = (-1/4)(-3) - 4 5 = 3/4 - 4 5 = 3 1/4 <-------- not true falsee